OpenStudy (anonymous):

Decide if the set {[0],[2],[4],[6],[8]} is in Z10 is a group under multiplication. If yes, is it abelian? If not, which conditions fail to hold?

OpenStudy (anonymous):

\[{[0],[2],[4],[6],[8]} \subseteq Z _{10}\]

terenzreignz (terenzreignz):

What exactly do the brackets mean?

terenzreignz (terenzreignz):

Well, see \[\huge \left\{0,2,4,6,8\right\}=\left<2\right>\subseteq \mathbb{Z}_{10}\] Under addition, this is simply a cyclic subgroup, but under multiplication...there is no multiplicative identity, so no.

OpenStudy (anonymous):

I am not sure about the brackets. That is how the example is written in my book.

OpenStudy (anonymous):

How do you know that 2 is not in there?

terenzreignz (terenzreignz):

Well, until further notice, I hope you don't mind that I left the brackets out...

terenzreignz (terenzreignz):

How do I know that 2 is not in there? 2 is there, all right...

OpenStudy (anonymous):

Yeah that is fine.

terenzreignz (terenzreignz):

Are you sure the question asks if it's a group under *multiplication*?

OpenStudy (anonymous):

Yes. It says decide if the set .... is a group under multiplication.

terenzreignz (terenzreignz):

Okay... a group needs four things... An associative binary operation Closure Existence of an identity Existence of inverses.... Let's check these... shall we?

terenzreignz (terenzreignz):

Is multiplication associative?

OpenStudy (anonymous):

Yes. I understand that.

OpenStudy (anonymous):

Yes multiplication is associative.

terenzreignz (terenzreignz):

Okay, when you multiply two elements of the set {0,2,4,6,8} is the product still in the set?

terenzreignz (terenzreignz):

Or we could just cut to the chase :)

OpenStudy (anonymous):

No.

terenzreignz (terenzreignz):

Yes, it is, actually closed under multiplication (in \(\large \mathbb{Z}_{10}\) )

terenzreignz (terenzreignz):

But more importantly, the existence of an identity element.

OpenStudy (anonymous):

\[4 * 6 = 24, \not \in (\mathbb{Z} _{10})\]. So how? I am not sure how to find the identity element

terenzreignz (terenzreignz):

No, see (this must be one of my catchphrases by now, along with "catch me so far?" LOL) Whenever you add or multiply in a set \(\large \mathbb{Z}_n\) you always take the answer in modulo n. So actually \[\huge 4\cdot 6 = 24(\mod 10 \ \ )=4\in\mathbb{Z}_{10}\]

OpenStudy (anonymous):

Okay I got you on that

terenzreignz (terenzreignz):

Which just made me realise... there IS an identity element after all!

OpenStudy (anonymous):

How do I know if something is an identity element?

terenzreignz (terenzreignz):

\[\huge 6\cdot 0 = 0\]\[\huge6\cdot2=12(\mod 10 \ \ )=2\]\[\huge 6\cdot 4=23(\mod 10 \ \ )=4\]\[\huge 6\cdot 6 = 36 \rightarrow 6\]\[\huge 6\cdot 8 =48\rightarrow 8\]

terenzreignz (terenzreignz):

If there's an element, that, when multiplied (?) to any other element, the result is that element.

terenzreignz (terenzreignz):

As you can see, 6 seems to satisfy that.

OpenStudy (anonymous):

Okay I got you.

terenzreignz (terenzreignz):

So far so good?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

So, apparently, it's the fourth one that presents a problem... You agree that 6 is the identity element, right?

OpenStudy (anonymous):

Yes I do.

terenzreignz (terenzreignz):

Well, to complete the circle, every element in the group must have an element you can multiply to it so that you get the identity 6 (its inverse) \[2\cdot 8 = 6\]\[4\cdot 4 = 6\]\[6\cdot 6 = 6\]\[8\cdot2=6\]

terenzreignz (terenzreignz):

Did I leave anything out?

OpenStudy (anonymous):

I have to go. I will be back later to get a better understanding of the inverse.

terenzreignz (terenzreignz):

Well, anyway, there simply is nothing you can multiply to 0 so that you get 6, no matter what modulo you're using. Hence, 0 has no inverse Hence, this is no group. Get it? Got it? Good :) ------------------------------------------------- Terence out

OpenStudy (anonymous):

So would this be abelian?

terenzreignz (terenzreignz):

It is not even a group.